OpenStudy (anonymous):

Verify the Identity. tan(x+pi/2)=-cot x

OpenStudy (anonymous):

do you know the formula for tan (A+B)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

use it. sub x for A and pi/2 for B

OpenStudy (anonymous):

remember that sin(pi/2) = 1 and cos(pi/2) = 0

OpenStudy (anonymous):

tan (a+b)=sin (a+b)/cos(a+b)

OpenStudy (anonymous):

?

terenzreignz (terenzreignz):

Yes, and don't forget also the identities for sin(a+b) and cos(a+b) Like sin(a+b) = sin(a)cos(b) + cos(a)sin(b) cos(a+b) = cos(a)cos(b) - sin(a)sin(b)

OpenStudy (anonymous):

i think i remember it as \[\frac{\tan A + \tan B}{1 + \tan A\tan B}\] or something like that

OpenStudy (anonymous):

maybe i'm wrong

terenzreignz (terenzreignz):

Can't, @abb50 because... What will you do about \[\Large \tan \frac\pi 2=\color{red}?\]

OpenStudy (anonymous):

interesting point

OpenStudy (anonymous):

I know that tan(x-pi/2)=cot but i dont know how to get to it

terenzreignz (terenzreignz):

Plug it in here... \[\Large \tan \left(x + \frac \pi 2\right)= \frac{\sin \left(x + \frac \pi 2\right)}{\cos \left(x + \frac \pi 2\right)}\]

OpenStudy (anonymous):

I still dont understand

OpenStudy (anonymous):

since a 99 SS user has come, I suppose I'll take my leave now

terenzreignz (terenzreignz):

No, by all means, carry on, @abb50 if you like :)

OpenStudy (anonymous):

Maybe you can try to convert \(-cot(x)\) to\(\frac{ -1 }{ \tan(x) }\), then use the \(\tan(a+b)\)formula to figure out what the left side is.

OpenStudy (jdoe0001):

$$ \Large \tan \left(x + \frac \pi 2\right)= \frac{\sin \left(x + \frac \pi 2\right)}{\cos \left(x + \frac \pi 2\right)} \implies \frac{cos(x)}{-sin(x)} $$

OpenStudy (anonymous):

@terenzreignz is not a 99 SS (at least not yet) he is a 96 SS which is great :)

terenzreignz (terenzreignz):

cross fingers :3

OpenStudy (anonymous):

i still dont understand

terenzreignz (terenzreignz):

anyway, @tpw1205 here's the deal :D I'll expand the numerator, you expand the denominator, ok? Here goes.. \[\Large \frac{\sin \left(x + \frac \pi 2\right)}{\cos \left(x + \frac \pi 2\right)}=\frac{\color{blue}{\sin(x)\cos\left(\frac \pi 2\right)+\cos(x)\sin\left(\frac \pi 2\right)}}{\cos\left(x+\frac \pi 2\right)}\]

OpenStudy (anonymous):

ok

terenzreignz (terenzreignz):

I used this identity: \[\Large \sin(\alpha+\beta) = \sin(\alpha)\cos(\beta) +\cos(\alpha)\sin(\beta)\] Use THIS identity with the denominator: \[\Large \cos(\alpha+\beta) = \cos(\alpha)\cos(\beta) -\sin(\alpha)\sin(\beta)\]

terenzreignz (terenzreignz):

Where \[\large \alpha = x\]\[\large \beta = \frac \pi 2\]

OpenStudy (anonymous):

thanks